The length of a rectangle is twice the width. The area is 8yd^2. Find the length and the width.
The length is ____yd
The width is _______yd

Answer by mananth(16949)

width be x
length = 2x
Area = L*W
x*2x =8
2x^2=8
x^2= 8/2
x^2=4
x= sqrt 4
x=2 the width
length = 2x = 2*2 =4